OpenStudy (anonymous):

To the nearest hundredth, what is the area of a circle with a circumference of 22 units ?

OpenStudy (mayankdevnani):

\[\huge circumference ~ of ~ circle=2 \pi r\] first, try to find the value of r where pi=22/7

OpenStudy (mayankdevnani):

\[\huge 22=2 \times \frac{22}{7} \times r\] find the value of r??? @atrabilious

OpenStudy (anonymous):

Hold on, (:

OpenStudy (jhannybean):

C = 2*pi*r C = 22 22/2 pi = r 11 pi = r A= pi*r^2 A = pi* (11 pi)^2 A = pi * 121/pi^2 A= ?

OpenStudy (mayankdevnani):

sure.. @atrabilious

OpenStudy (anonymous):

Oh no my answer isn't right at all.

OpenStudy (jhannybean):

lmao \[\huge 22 = 2 \pi r \therefore r = 22/2\pi = 11 \pi\]

OpenStudy (anonymous):

38.52 ?

OpenStudy (mayankdevnani):

ok...i give you hint: \[\huge r=\frac{22 \times 7}{2 \times 22}\] can you solve it?? @atrabilious

OpenStudy (anonymous):

Wtf sorry for not seeing that till right now @Jhannybean

OpenStudy (jhannybean):

why are you multiplying by 7? lol

OpenStudy (anonymous):

34.56 (:

OpenStudy (anonymous):

units

OpenStudy (mayankdevnani):

 ok.. @Jhannybean